OpenStudy (anonymous):

(21/25)(10/117)(27/14) multiply. give answer in lowest term. I am terrible at finding the LCD, is that where I need to start? or Multiply and then find simplest form?

OpenStudy (anonymous):

all u gotta do is just put it in ur calculator..or r u supposed to do it by hand?

OpenStudy (anonymous):

Just multiply them out, then reduce the answer (is what I'd do, at least)

OpenStudy (amistre64):

21.10.27 --------- 25.117.14 3. 1 . 3 ------- = 6 / 65 5.13. 1

OpenStudy (anonymous):

did you factor out the numbers?

OpenStudy (amistre64):

i tried.... let me see if i can step thru that to check meself

OpenStudy (amistre64):

21.10.27 --------- 25.117.14 21.2.27 --------- 5.117.14 3.2.27 ------- 5.117.2 3.2.9.3 ------- 5.9.13.2 3.3 ---- = 6/65 5.13

OpenStudy (anonymous):

21.10.27 --------- 25.117.14 21.1.27 ------- 5.117.7 3.1.9 ----- 5.39.1 1.1.9 ----- 5.13.1 =9/65

OpenStudy (anonymous):

thank you both!

OpenStudy (amistre64):

9/65 is a better fit :)

OpenStudy (amistre64):

LOL!!... o forgot how to multiply 3.3 LOL!!!!

OpenStudy (anonymous):

i noticed that when i checked through it, haha. but thanks for all the work as well!

OpenStudy (amistre64):

the embarassing part is.... i did it twice lol

OpenStudy (anonymous):

:) it happens!

OpenStudy (anonymous):

third times a charm ;) damn public libraries and their sticky keys!